OpenStudy (sidsiddhartha):

All the values of the multivalued complex function \[\Large 1^{i}\] are-- (1)purely imaginary (2)real and non negetive (3)on the unit circle (4)equal in real and imaginary parts

OpenStudy (sidsiddhartha):

i'm going for option B anyone agrees with me?

OpenStudy (anonymous):

I would agree, since \(1^i=1\).

OpenStudy (sidsiddhartha):

yeah thanks \[1=e^{i(2n \pi)}\rightarrow 1^i=e^{-2n \pi}\] right ?

OpenStudy (kainui):

Well if you look at this: \[\LARGE e^{i \theta} = \cos \theta + i \sin \theta\] We can quickly see that if we plug in theta = 0 we get: \[\LARGE e^{i 0} = \cos 0+ i \sin 0\]\[\LARGE (e^0)^i = 1+i0\]\[\LARGE 1^i=1\] It seems to me that both 2 and 3 are correct since the value 1 lies on the unit circle and is both real and nonnegative.